OpenStudy (anonymous):

If h(x) = −2x − 10, find h(−4). −2 −18 −3 −16

OpenStudy (australopithecus):

h(-4) is just the function h(x), with -4 subbed in as x

OpenStudy (anonymous):

plug in -4 for x =-2(-4) - 10 = 8 - 10 = -2

OpenStudy (australopithecus):

so replace x by -4 to solve